OpenStudy (anonymous):

The mini pizza has a 6-inch diameter, and the medium pizza has a diameter twice that of the mini pizza. How many more square inches of pizza are there in the medium than in the mini pizza? Round your answer to the nearest tenth.

OpenStudy (anonymous):

A = pi*r^2 2r = d 1/2d = r Mini Pizza] d = 6 r =3 A = pi * 3^2 = 9pi Medium Pizza] d = 6*2 = 12 r = 6 A = pi * 6^2 = 36pi Difference: 36pi - 9pi = 27pi pi = 3.14159 27 * 3.14159 = 75.39816 Rounded to tenths: 75.4 square inches larger

OpenStudy (anonymous):

thank you very much(: